# Discrete Heat Kernels on Simplicial Complexes and Its Application to Functional Brain Networks

### v1

@@ 2.3.4 Modularity and Random Graphs
The concept of modularity of networks or graphs is associated with the strength of the division of networks or graphs into modules. Modularity is therefore a quality index for clustering graphs [37].
Consider a graph $\mathcal{X}=(V,E)$ with $|V|=p$ and let $\mathcal{C}=\{C_{1},...,C_{K}\}$ be a partition of $V$. If $C_{k}\neq\emptyset$ $\forall k\in\{1,...,K\}$, then $\mathcal{C}$ is known as a clustering of $\mathcal{X}$. It can be observed that $\mathcal{C}$ will be trivial if either $K$ is 1 or $p$. We can view each $C_{k}$ as an induced subgraph $\mathcal{X}_{K}=\mathcal{X}\left[C_{k}\right]=(C_{k},E(C_{k}))$, where $E(C_{k})=\{\{v_{i},v_{j}\}\in E:v_{i},v_{j}\in C_{k}\}$. The strength of the modular structure defined by the partition $\mathcal{C}$ can be quantified via the modularity measure [37]
@@
where $d(v_{i})$ is the degree of node $v_{i}$. The objective is always to maximize $Q$ which involves maximizing the number of intramodules edges and minimizing the number of intermodule edges. The maximum value of $Q$ is 1, and values close to that indicates strong modular structure. However, empirical modularity values of graphs falls between $0.3$ and $0.7$ [37].
@@ 2.3.5 Weighted Random Modular Graphs
Various methods exist for the construction of weighted random modular graphs. For example, the assignment of nodes to modules can be based on the module size distribution, and edge connection modeled after the Havel-Hakimi algorithm [37]. Others also propose modeling edge weights based of Gaussian mixture models (GMM) or the use of GMM coupled with variational inference [23]. All these methods share a common theme of devising a reasonable probability distribution to model edge weights. In what follows, we will describe an intuitive and simple process for generating random graphs and assigning random weights to generated edges. This was inspired by [46].
@@ 2.4 The Graph Laplacian
We will now briefly introduce the concept of orientation of a graph that will helps us prove certain properties of $\mathcal{L}_{0}$.
Consider the graph $\mathcal{X}=(V,E)$, undirected, an orientation of $\mathcal{X}$ is a mapping $f:E\xrightarrow{}V\times V$ assigning a source(s) and target(t) to every edge in $E$, i.e., for every edge $\{u,v\}\in E$, either f(u,v)=(u,v)f({u,v})=(u,v) or f(u,v)=(v,u)f({u,v})=(v,u). The oriented graph $\mathcal{X}^{o}$ derived from $\mathcal{X}$ by applying the orientation $f$ is $\mathcal{X}^{0}=(X,E^{0})$, $E^{0}=f(E)$.
Denote the incidence matrix associated with $\mathcal{X}^{o}$ by $B^{o}$ which is a $|V|\times|E|$ with entries
@@ 2.5 Hodge Laplacian
in the matrix form, which is mainly used for actual computation. If there is no ambiguity, we will intermix the boundary operator and its matrix form. Here $\partial_{k}^{*}$ is the adjoint of $\partial_{k}$. It’s straightforward to check that, $\mathcal{L}_{k}$ is a $|\mathcal{K}_{k}|\times|\mathcal{K}_{k}|$ matrix, where $|\mathcal{K}_{k}|$ is the cardinality of k-simplices in $\mathcal{K}_{k}$. The Hodge Laplacian has been very useful not only in establishing various theoretical results but also in various applied works [44, 35].
@@
Fig. 3: A schematic of a 1-skeleton representation of a simplicial complex $\mathcal{K}$. An oriented simplicial complex $\mathcal{K}$ (left) with each filled in triangle denoting a 2-simplex. A graph corresponding to the 1-skeleton representation of $\mathcal{K}$.
@@
Figure 3 provides an illustration of such a mapping. We can check that the matrix representation $\mathbb{B}_{1}\mathbb{B}_{1}^{T}$ will have entries corresponding to the unweighted version in Proposition 3. An alternative construction of the Laplacain $\mathcal{L}_{k}$ for $k>0$ is as follows. Two simplices $\sigma_{i}$, $\sigma_{j}$ are said to be upper adjacent if both are faces of some $k+1$-simplex in $\mathcal{K}$. Denote this adjacency by $\sigma_{i}\wedge\sigma_{j}$. The upper degree of a k-simplex $(deg_{u}(\sigma))$ is the number of $k+1$-simplices in $\mathcal{K}$ of which $\sigma$ is a face. The above generalizes the concept of degree and adjacency of graphs, and note that analogous definitions holds for the lower degree and upper adjacency. It can be shown that [35]
@@ 2.5.1 Cycles
We now extend the definition of a cycle under 1-skeleton to any simplicial complex of arbitrary order. In previous section, we defined the $k$-th homology group as ${\cal H}_{k}={\cal Z}_{k}/{\cal B}_{k}$ where ${\cal Z}_{k}$ is the $k$-cycle group and ${\cal B}_{k}$ is the $(k+1)$-boundary group with ${\cal B}_{k}\subset{\cal Z}_{k}$. The rank of the homology groups gives the Betti numbers $\beta_{k}=rank({\cal H}_{k})$ [28, 49]. Since the $k$-th homology group contains k-dimensional cycles which differ by only homological cycles, the $\beta_{k}$ counts the number of k-cycles. The basis associated with ${\cal H}_{0}$, ${\cal H}_{1}$ and ${\cal H}_{2}$ corresponds to connected components, holes and cavities respectively.
We will now motivate the concept of persistence of k-cycles through persistent homology. Let $\cal K$} be a simplicial complex, and analogous to graph filtration, consider a filtration $\{{\cal K}_{\epsilon_{k}}\}_{\epsilon_{k}\in\mathbb{R}}$ where for $\epsilon_{k_{1}}\leq\epsilon_{k_{2}}$, we have that ${\cal K}_{\epsilon_{k_{1}}}\subset{\cal K}_{\epsilon_{k_{2}}}\subset{\cal K}$. The filtration of $\cal K$} induces a homomorphism between the homology groups, i.e. ${\cal H}_{k}\left({\cal K}_{\epsilon_{k_{1}}}\right)\xrightarrow[]{}{\cal H}_{k}\left({\cal K}_{\epsilon_{k_{2}}}\right)$, which allows us to track the birth of homology classes over the span of the filtration to their death point. Persistence modules, which uniquely decompose into persistence intervals [53] is a tool for tracking this evolution. Let ${\cal P}_{k}$ denote the persistence module, and for a $k$-cycle ${\cal C}\in{\cal P}_{k}$, denote the point it first appeared (birth) as $b({\cal C})$ and the point it disappeared (death) as $d({\cal C})$. Then the lifespan of this cycle can be represented as $\left[b({\cal C}),d({\cal C})\right)$, the persistence interval. A finite collection of the persistence interval $\left[b({\cal C}),d({\cal C})\right)$ can be summarized in the form of a barcode, which can intuitively be thought of as the persistence analogue of Betti numbers.
@@ 2.5.2 Diffusion on Simplicial Complexes
where $f_{j}=f^{\top}\psi_{j}$ is the Fourier coefficient with respect to $\psi_{j}$ [30, 42].
We can generalize this result to k-skeletons. Assume $g(t)$ to be functional data with some parameter $t$ at a generic simplex $\sigma$ (a collection of k-simplices). It is expected such functional data is highly noisy and requires denoising, which will be done through the heat equation. Before we formulate the generalized model, we state some important properties associated with the Hodge Laplacian $\mathcal{L}_{k}$ for a general $k$. There are quite a number of important properties of the eigenvectors and eigenvalues of $\mathcal{L}_{K}$ [6]. One important property is the symmetricity of $\mathcal{L}_{k}$, which also implies its diagonalizable. Ordinarily, been symmetric is the property of operators and not just the associated matrix, but as indicated earlier we will intermix the boundary operator and the matrix form. Henceforth, properties defined will be assumed to be on the matrix form of $\mathcal{L}_{k}$.
@@ 2.6 Heat kernel over simplices
where $t$ is the kernel bandwidth. With the exception of a few geometries, the heat kernel $K_{t}$ does not admit an explicit analytic form since the eigenfunctions $\psi_{j}$ are not often given analytically [12]. A proposed workaround is to study the small bandwidth behavior via the parametrix expansion, the asymptotic expansion of heat kernel [43]. When the geodesic distance $d(\sigma_{1},\sigma_{2})$ between simplices $\sigma_{1}$ and $\sigma_{2}$ is small enough, we have the following parametrix expansion [12]
@@
then the metric space is flat, in which case the dimension of the metric does not exceed a specified Euclidean kk+11-space [34]. The metric space can then be embedded in the Euclidean space, and the heat kernel becomes a Gaussian kernel [12, 34]. If we convolve $K_{t}$ with the initial functional data, we can algebraically show that the convolution is solution to (25) [12, 30]:
@@ 2.7 Simulation Study
Fig. 4: An illustration of the simulated and smoothed data. On the far-left is the collection of 150 0-simplices with the color gradient indicating the initial concentration at each of the nodes. We notice the color spectrum is widely varying as expected. In the middle and far-right is the resulting 0-simplices after smoothing via the heat diffusion equation. Its apparent that there is a greater uniformity in the distribution of the color gradient compared to the unsmoothed one. This signals the denoising of the initial functional data defined on the nodes.
@@
Fig. 5: This illustrate the diffusion process on 1-simplices. On the far-left is the collection of 1-simplices with the initial unsmoothed functional data. Again the color gradient is quite varying. The middle and far-right graph depicts the output when the data is denoised. We observe that color gradient across the 1-simplices is relatively uniform compared to the unsmoothed data.
@@ 2.8 Representation of k-cycles
Similar to the realization in [2], even k-simplices that do not form part of the k-cycle are considered in (29). The way to solve this problem for the special case of $k=1$ is to sequentially construct the 1-cycles drawing from the theory of spanning trees [2, 10]. This approach is not feasible for the case where $k>1$. But we can have an analogous construction that retains the core property that the entries of the eigenvectors will be zero for k-simplices that are not part of the cycle.
@@
Fig. 6: Top: Group-mean functional connectivity of users on a glass brain. Relatively few negative connections were observed. Bottom: Group-mean functional connectivity of non-users on a glass brain. Relatively few negative connections were observed. The dense representation of the connectivities gives no visibly discernible pattern.
@@ 3.1 Data and Preprocessing
The data analyzed is the Addiction Connectome Preprocessed Initiative (ACPI) distribution of the Multimodal Treatment of ADHD (MTA) resting-state fMRI resource released through FCP/INDI. This is a derivative, preprocessed release: raw scans were curated by the project and distributed both as standardized 4D volumes and as parcellated region–wise time series. The parcellated time series in the Automated Anatomical Labeling (AAL) atlas was used for this analysis [50]. The data is in two groups based on marijuana use status, group that regularly use cannabis (users) and another group that did not regularly use cannabis (non-users).
@@
Fig. 7: Top: he functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.05$ of the mean connectivity for users. Observe that spurious isolated connections are attenuated (connectivity reduced) while coherent bundles are enhanced. Bottom: The functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.05$ of the mean connectivity for non-users. Observe that spurious isolated connections are attenuated (connectivity reduced) while coherent bundles are enhanced.
@@ 3.2 Results
The brain networks in Figure 6 shows the unsmoothed group–mean connectivity for users and non–users respectively. The matrices display the expected within–lobe clustering and interhemispheric homologous links. Observe the dense representation of the connectivities with no visibly discernible pattern. Neuroimaging studies have shown how cannabis affects brain connectivity patterns. Studies examining functional brain connectomes have revealed that both acute and chronic cannabis exposure alter resting-state connectivity, with particular emphasis on networks involved in executive control and self-referential processing[41]. THC has been shown to disrupt the DMN, with the posterior cingulate cortex (PCC) being a key region involved in the subjective experience of THC intoxication [51]. The results in the next section demonstrate that heat kernel smoothing enhance these specific regional patterns without imposing artificial boundaries, which could be very useful in studying such pharmacologically-induced changes.
@@
Fig. 8: Top: The functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.1$ of the mean connectivity for users. Observe that spurious isolated connections are more attenuated compared to the case where the bandwidth is $t=0.05$. This suppresses isolated noisy connections and sharpens coherent patterns without imposing modular structure or hard thresholds. Bottom: The functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.1$ of the mean connectivity for non-users. Observe that spurious isolated connections are more attenuated compared to the case where the bandwidth is $t=0.05$. This suppresses isolated noisy connections and sharpens coherent patterns without imposing modular structure or hard thresholds.

### v2

@@ 2.3.4 Modularity and Random Graphs
The concept of modularity of networks or graphs is associated with the strength of the division of networks or graphs into modules. Modularity is therefore a quality index for clustering graphs [newman2006modularity].
Consider a graph $\mathcal{X}=(V,E)$ with $|V|=p$ and let $\mathcal{C}=\{C_{1},...,C_{K}\}$ be a partition of $V$. If $C_{k}\neq\emptyset$ $\forall k\in\{1,...,K\}$, then $\mathcal{C}$ is known as a clustering of $\mathcal{X}$. It can be observed that $\mathcal{C}$ will be trivial if either $K$ is 1 or $p$. We can view each $C_{k}$ as an induced subgraph $\mathcal{X}_{K}=\mathcal{X}\left[C_{k}\right]=(C_{k},E(C_{k}))$, where $E(C_{k})=\{\{v_{i},v_{j}\}\in E:v_{i},v_{j}\in C_{k}\}$. The strength of the modular structure defined by the partition $\mathcal{C}$ can be quantified via the modularity measure [newman2006modularity]
@@
where $d(v_{i})$ is the degree of node $v_{i}$. The objective is always to maximize $Q$ which involves maximizing the number of intramodules edges and minimizing the number of intermodule edges. The maximum value of $Q$ is 1, and values close to that indicates strong modular structure. However, empirical modularity values of graphs falls between $0.3$ and $0.7$ [newman2006modularity].
@@ 2.3.5 Weighted Random Modular Graphs
Various methods exist for the construction of weighted random modular graphs. For example, the assignment of nodes to modules can be based on the module size distribution, and edge connection modeled after the Havel-Hakimi algorithm [newman2006modularity]. Others also propose modeling edge weights based of Gaussian mixture models (GMM) or the use of GMM coupled with variational inference [garlaschelli2009weighted]. All these methods share a common theme of devising a reasonable probability distribution to model edge weights. In what follows, we will describe an intuitive and simple process for generating random graphs and assigning random weights to generated edges. This was inspired by [songdechakraiwut2020topological].
@@ 2.4 The Graph Laplacian
We will now briefly introduce the concept of orientation of a graph that will helps us prove certain properties of $\mathcal{L}_{0}$. Consider the graph $\mathcal{X}=(V,E)$, undirected, an orientation of $\mathcal{X}$ is a mapping $f:E\xrightarrow{}V\times V$ assigning a source(s) and target(t) to every edge in $E$, i.e., for every edge $\{u,v\}\in E$, either f(u,v)=(u,v)f({u,v})=(u,v) or f(u,v)=(v,u)f({u,v})=(v,u). The oriented graph $\mathcal{X}^{o}$ derived from $\mathcal{X}$ by applying the orientation $f$ is $\mathcal{X}^{0}=(X,E^{0})$, $E^{0}=f(E)$. Denote the incidence matrix associated with $\mathcal{X}^{o}$ by $B^{o}$ which is a $|V|\times|E|$ with entries
@@ 2.5 Hodge Laplacian
in the matrix form, which is mainly used for actual computation. If there is no ambiguity, we will intermix the boundary operator and its matrix form. Here $\partial_{k}^{*}$ is the adjoint of $\partial_{k}$. It’s straightforward to check that, $\mathcal{L}_{k}$ is a $|\mathcal{K}_{k}|\times|\mathcal{K}_{k}|$ matrix, where $|\mathcal{K}_{k}|$ is the cardinality of k-simplices in $\mathcal{K}_{k}$. The Hodge Laplacian has been very useful not only in establishing various theoretical results but also in various applied works [schaub2020random, muhammad2006control].
@@
Figure 3: A schematic of a 1-skeleton representation of a simplicial complex $\mathcal{K}$. An oriented simplicial complex $\mathcal{K}$ (left) with each filled in triangle denoting a 2-simplex. A graph corresponding to the 1-skeleton representation of $\mathcal{K}$.
@@
Figure 3 provides an illustration of such a mapping. We can check that the matrix representation $\mathbb{B}_{1}\mathbb{B}_{1}^{T}$ will have entries corresponding to the unweighted version in Proposition 3. An alternative construction of the Laplacain $\mathcal{L}_{k}$ for $k>0$ is as follows. Two simplices $\sigma_{i}$, $\sigma_{j}$ are said to be upper adjacent if both are faces of some $k+1$-simplex in $\mathcal{K}$. Denote this adjacency by $\sigma_{i}\wedge\sigma_{j}$. The upper degree of a k-simplex $(deg_{u}(\sigma))$ is the number of $k+1$-simplices in $\mathcal{K}$ of which $\sigma$ is a face. The above generalizes the concept of degree and adjacency of graphs, and note that analogous definitions holds for the lower degree and upper adjacency. It can be shown that [muhammad2006control]
@@ 2.5.1 Cycles
We now extend the definition of a cycle under 1-skeleton to any simplicial complex of arbitrary order. In previous section, we defined the $k$-th homology group as ${\cal H}_{k}={\cal Z}_{k}/{\cal B}_{k}$ where ${\cal Z}_{k}$ is the $k$-cycle group and ${\cal B}_{k}$ is the $(k+1)$-boundary group with ${\cal B}_{k}\subset{\cal Z}_{k}$. The rank of the homology groups gives the Betti numbers $\beta_{k}=rank({\cal H}_{k})$ [hatcher2002algebraic, topaz2015topological]. Since the $k$-th homology group contains k-dimensional cycles which differ by only homological cycles, the $\beta_{k}$ counts the number of k-cycles. The basis associated with ${\cal H}_{0}$, ${\cal H}_{1}$ and ${\cal H}_{2}$ corresponds to connected components, holes and cavities respectively. We will now motivate the concept of persistence of k-cycles through persistent homology. Let $\cal K$} be a simplicial complex, and analogous to graph filtration, consider a filtration $\{{\cal K}_{\epsilon_{k}}\}_{\epsilon_{k}\in\mathbb{R}}$ where for $\epsilon_{k_{1}}\leq\epsilon_{k_{2}}$, we have that ${\cal K}_{\epsilon_{k_{1}}}\subset{\cal K}_{\epsilon_{k_{2}}}\subset{\cal K}$. The filtration of $\cal K$} induces a homomorphism between the homology groups, i.e. ${\cal H}_{k}\left({\cal K}_{\epsilon_{k_{1}}}\right)\xrightarrow[]{}{\cal H}_{k}\left({\cal K}_{\epsilon_{k_{2}}}\right)$, which allows us to track the birth of homology classes over the span of the filtration to their death point. Persistence modules, which uniquely decompose into persistence intervals [zomorodian2005computing] is a tool for tracking this evolution. Let ${\cal P}_{k}$ denote the persistence module, and for a $k$-cycle ${\cal C}\in{\cal P}_{k}$, denote the point it first appeared (birth) as $b({\cal C})$ and the point it disappeared (death) as $d({\cal C})$. Then the lifespan of this cycle can be represented as $\left[b({\cal C}),d({\cal C})\right)$, the persistence interval. A finite collection of the persistence interval $\left[b({\cal C}),d({\cal C})\right)$ can be summarized in the form of a barcode, which can intuitively be thought of as the persistence analogue of Betti numbers.
@@ 2.5.2 Diffusion on Simplicial Complexes
where $f_{j}=f^{\top}\psi_{j}$ is the Fourier coefficient with respect to $\psi_{j}$ [huang2019fast, RICAUD2019474].
We can generalize this result to k-skeletons. Assume $g(t)$ to be functional data with some parameter $t$ at a generic simplex $\sigma$ (a collection of k-simplices). It is expected such functional data is highly noisy and requires denoising, which will be done through the heat equation. Before we formulate the generalized model, we state some important properties associated with the Hodge Laplacian $\mathcal{L}_{k}$ for a general $k$. There are quite a number of important properties of the eigenvectors and eigenvalues of $\mathcal{L}_{K}$ [barbarossa2018learning]. One important property is the symmetricity of $\mathcal{L}_{k}$, which also implies its diagonalizable. Ordinarily, been symmetric is the property of operators and not just the associated matrix, but as indicated earlier we will intermix the boundary operator and the matrix form. Henceforth, properties defined will be assumed to be on the matrix form of $\mathcal{L}_{k}$.
@@ 2.6 Heat kernel over simplices
where $t$ is the kernel bandwidth. With the exception of a few geometries, the heat kernel $K_{t}$ does not admit an explicit analytic form since the eigenfunctions $\psi_{j}$ are not often given analytically [chung2015unified]. A proposed workaround is to study the small bandwidth behavior via the parametrix expansion, the asymptotic expansion of heat kernel [rosenberg1997laplacian]. When the geodesic distance $d(\sigma_{1},\sigma_{2})$ between simplices $\sigma_{1}$ and $\sigma_{2}$ is small enough, we have the following parametrix expansion [chung2015unified]
@@
then the metric space is flat, in which case the dimension of the metric does not exceed a specified Euclidean kk+11-space [morgan1974embedding]. The metric space can then be embedded in the Euclidean space, and the heat kernel becomes a Gaussian kernel [chung2015unified, morgan1974embedding]. If we convolve $K_{t}$ with the initial functional data, we can algebraically show that the convolution is solution to (25) [chung2015unified, huang2019fast]:
@@ 2.7 Simulation Study
Figure 4: An illustration of the simulated and smoothed data. On the far-left is the collection of 150 0-simplices with the color gradient indicating the initial concentration at each of the nodes. We notice the color spectrum is widely varying as expected. In the middle and far-right is the resulting 0-simplices after smoothing via the heat diffusion equation. Its apparent that there is a greater uniformity in the distribution of the color gradient compared to the unsmoothed one. This signals the denoising of the initial functional data defined on the nodes.
@@
Figure 5: This illustrate the diffusion process on 1-simplices. On the far-left is the collection of 1-simplices with the initial unsmoothed functional data. Again the color gradient is quite varying. The middle and far-right graph depicts the output when the data is denoised. We observe that color gradient across the 1-simplices is relatively uniform compared to the unsmoothed data.
@@ 2.8 Representation of k-cycles
Similar to the realization in [anand2021hodge], even k-simplices that do not form part of the k-cycle are considered in (29). The way to solve this problem for the special case of $k=1$ is to sequentially construct the 1-cycles drawing from the theory of spanning trees [anand2021hodge, busaryev2012annotating]. This approach is not feasible for the case where $k>1$. But we can have an analogous construction that retains the core property that the entries of the eigenvectors will be zero for k-simplices that are not part of the cycle.
@@
Figure 6: Top: Group-mean functional connectivity of users on a glass brain. Relatively few negative connections were observed. Bottom: Group-mean functional connectivity of non-users on a glass brain. Relatively few negative connections were observed. The dense representation of the connectivities gives no visibly discernible pattern.
@@ 3.1 Data and Preprocessing
The data analyzed is the Addiction Connectome Preprocessed Initiative (ACPI) distribution of the Multimodal Treatment of ADHD (MTA) resting-state fMRI resource released through FCP/INDI. This is a derivative, preprocessed release: raw scans were curated by the project and distributed both as standardized 4D volumes and as parcellated region–wise time series. The parcellated time series in the Automated Anatomical Labeling (AAL) atlas was used for this analysis [tzourio2002automated]. The data is in two groups based on marijuana use status, group that regularly use cannabis (users) and another group that did not regularly use cannabis (non-users).
@@
Figure 7: Top: he functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.05$ of the mean connectivity for users. Observe that spurious isolated connections are attenuated (connectivity reduced) while coherent bundles are enhanced. Bottom: The functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.05$ of the mean connectivity for non-users. Observe that spurious isolated connections are attenuated (connectivity reduced) while coherent bundles are enhanced.
@@ 3.2 Results
The brain networks in Figure 6 shows the unsmoothed group–mean connectivity for users and non–users respectively. The matrices display the expected within–lobe clustering and interhemispheric homologous links. Observe the dense representation of the connectivities with no visibly discernible pattern. Neuroimaging studies have shown how cannabis affects brain connectivity patterns. Studies examining functional brain connectomes have revealed that both acute and chronic cannabis exposure alter resting-state connectivity, with particular emphasis on networks involved in executive control and self-referential processing[ramaekers2022functional]. THC has been shown to disrupt the DMN, with the posterior cingulate cortex (PCC) being a key region involved in the subjective experience of THC intoxication [wall2019dissociable]. The results in the next section demonstrate that heat kernel smoothing enhance these specific regional patterns without imposing artificial boundaries, which could be very useful in studying such pharmacologically-induced changes.
@@
Figure 8: Top: The functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.1$ of the mean connectivity for users. Observe that spurious isolated connections are more attenuated compared to the case where the bandwidth is $t=0.05$. This suppresses isolated noisy connections and sharpens coherent patterns without imposing modular structure or hard thresholds. Bottom: The functionality connectivity after smoothing along the 1-simplices (edges) with bandwidth $t=0.1$ of the mean connectivity for non-users. Observe that spurious isolated connections are more attenuated compared to the case where the bandwidth is $t=0.05$. This suppresses isolated noisy connections and sharpens coherent patterns without imposing modular structure or hard thresholds.